List of Publications

[1] On the approach of Holmes and Sanders to the Melnikov procedure in the method of averaging, J. Appl. Math. Phys. (ZAMP), Vol. 34 (1983), pp. 948-952.

[2] Some geometric and numerical methods for perturbed integrable systems, Diss. ETH No. 8456, (1988).

[3] Transversal homoclinic points and hyperbolic sets for non-autonomous maps I, J. Appl. Math. (ZAMP), Vol. 39 (1988), pp. 518-549.

[4] Transversal homoclinic points and hyperbolic sets for non-autonomous maps II, J. Appl. Math. (ZAMP), Vol. 39 (1988), pp. 783-812.

[5] On reversible and canonical integration methods, SAM-Report 88-05, ETH-Zürich (1988).

[6] with U. Kirchgraber: On the definition of chaos, Z. angew. Math. Mech. (ZAMM), Vol. 69 (1989), pp. 175-185.

[7] with K. Palmer: Chaos in almost periodic systems, J. Appl. Math. (ZAMP), Vol. 40 (1989), pp. 592-602.

[8] with U. Kirchgraber: Chaotic behaviour in simple dynamical systems, SIAM Review, Vol. 32 (1990), pp. 424-452.

[9] with K. Nipp: Invariant curves for variable step size integrators, BIT, Vol. 31 (1991), pp. 169-180.

[10] with U. Kirchgraber, F. Lasagni, K. Nipp: On the application of invariant manifold theory, in particular to numerical analysis, Int. Series Num. Math., Vol. 97 (1991), pp. 189-197.

[11] with K. Nipp: Lineare Algebra, eine Einführung für Ingenieure unter besonderer Berücksichtigung numerischer Aspekte, Vdf Verlag der Fachvereine, Zürich (1991), 2. Auflage (1992), 3. Auflage (1994).

[12] with K. Nipp: Attractive invariant manifolds for maps: Existence, smoothness and continuous dependence on the map, SAM-Report 92-11, ETH-Zürich (1992).

[13] General linear methods: connection to one step methods and invariant curves, Numer. Math., Vol. 64 (1993), pp. 395-407.

[14] Geometric Behaviour of Numerical Integration Methods for ODEs, Habilitationsschrift ETH–Zürich (1994).

[15] with Ch. Lubich und K. Nipp: Runge-Kutta solutions of stiff differential equations near stationary points, SIAM J. Numer. Anal. Vol. 32 (1995), pp. 1296-1307.

[16] with K. Nipp: Invariant manifolds of numerical integration schemes applied to stiff systems of singular perturbation type - Part I: RK-methods, Numer. Math. Vol. 70 (1995), pp. 245-257. (Teil der Habilitationsschrift [14]).

[17] Variable steps for reversible integration methods, Computing Vol. 55 (1995), pp. 1-22. (Teil der Habilitationsschrift [14]).

[18] with M. Adelmeyer: Iteration einer n-Ecksabbildung, Elemente d. Math. 51 (1996), pp. 1-16.

[19] with K. Nipp: Invariant manifolds of numerical integration schemes applied to stiff systems of singular perturbation type - Part II: multistep methods, Numer. Math., Vol. 74 (1996), pp. 305-323.

[20] with E. Hairer: Reversible long-term integration with variable step size, SIAM J. Sci. Comput. 18 (1997), pp. 257–269.

[21] On the qualitative behaviour of symplectic integrators, Part I: Perturbed linear systems, Numer. Math. 77 (1997), pp. 535-547. Part II: Integrable systems, J. Math. Anal. Appl. 217 (1998), pp. 501-520. Part III: Perturbed integrable systems, J. Math. Anal. Appl. 217 (1998), pp. 521-545.

[22] with K.J. Palmer Rigorous verification of chaotic behaviour of maps using validated shadowing, Nonlinearity 12 (1999), pp. 1683-1698.

[23] with Urs Kirchgraber, Ulrich Manz Rigorous proof of chaotic behaviour in a dumbbell satellite model, JMAA 251 (2000),pp. 897–911.

[24] with U. Kirchgraber Verification of chaotic behaviour in the planar restricted three body problem, Appl.Num.Math. 39 (2001), pp. 415–433.

[25] with U. Kirchgraber Possible chaotic motion of comets in the Sun–Jupiter system — a computer–assisted approach based on shadowing, Nonlinearity 17 (2004), pp. 1–20.